### v1

# Customize-It-3D: High-Quality 3D Creation from A Single Image Using Subject-Specific Knowledge Prior

Abstract
In this paper, we present a novel two-stage approach that fully utilizes the information provided by the reference image to establish a customized knowledge prior for image-to-3D generation. While previous approaches primarily rely on a general diffusion prior, which struggles to yield consistent results with the reference image, we propose a subject-specific and multi-modal diffusion model. This model not only aids NeRF optimization by considering the shading mode for improved geometry but also enhances texture from the coarse results to achieve superior refinement. Both aspects contribute to faithfully aligning the 3D content with the subject. Extensive experiments showcase the superiority of our method, Customize-It-3D, outperforming previous works by a substantial margin. It produces faithful 360-degree reconstructions with impressive visual quality, making it well-suited for various applications, including text-to-3D creation. Our project page is: https://nnanhuang.github.io/projects/customize-it-3d/
Figure 1: Customize-It-3D has the capability to generate high-fidelity 3D content from just a single image. We present the normal map and novel-view renderings for general objects, highlighting the consistency texture and geometry with remarkable quality at novel views.
1 Introduction
3D creation from a single image remains a profoundly intricate challenge, due to its inherently ill-posed nature, arising from the significant information gap between a single 2D image and the comprehensive 3D spatial and textural properties of the depicted object. Despite its great challenge, recent works [59, 46, 38, 68] have achieved notable results in image-to-3D creation. They draw inspiration from text-to-3D methods [45, 64] which utilize score distillation sampling (SDS) loss to optimize a neural radiance field (NeRF). This SDS loss serves as a guiding principle for novel view synthesis, emulating the way humans utilize priors to infer 3D shape and texture from a single image. While these methods tend to produce promising results, they still exhibit noticeable inconsistencies when compared with the reference view, either in context of the geometry (i.e., multi-face issue) or from the texture perspective (novel views often lack fidelity, presenting smooth and disruptive perticulars).
When crafting intricate 3D models of a specific object, a professional artist indeed need not be equipped with an exhaustive understanding of all the objects in the natural world. Rather, the essential requirement lies in possessing specialized, profound knowledge pertaining to that specific subject at hand. This targeted expertise shall be more instrumental in the meticulous creation of 3D assets. Despite incorporating several adjustments such as CLIP loss and textual inversion considered in [59, 46], the supervision of novel views still relies on a general text-to-image model [49, 51, 53, 3, 18], which has challenge in producing consistent results based solely on text, given the inherent difficulty in crafting textual descriptions to encompass every detail. Consequently, these approaches struggle to maintain consistency and fall short in reconstructing high-fidelity 3D objects. This realization has inspired us to explore strategies for cultivating subject-specific knowledge prior, aiming to achieve high-quality image-to-3D creation.
We propose Customize-It-3D, a novel approach that fully unleashes the geometry/appearance information provided by the reference image to promote customized knowledge prior for 3D generation. Customize-It-3D, like the previous methods mentioned above, adopts a two-stage framework in a coarse-to-fine manner. However, it is guided with subject-specific knowledge prior.
In the first stage, we optimize a neural radiance field to learn an implicit volume representation. The reference view is directly supervised through rendering loss, specifically pixel differences. For novel views, unlike conventional approaches employing general T2I models for SDS loss, we use a subject-specific T2I model. There has been remarkable success in subject-driven image generation [52, 29]. They typically demand at least 3 to 5 images capturing the same subject with varying contexts in order to model the key visual features of the subject. However, our scenario presents a challenge as we possess only a solitary image, which may potentially misguide the model to a trivial solution (mode collapse). To surmount this limitation, we propose to exploit multiple modalities extracted from the reference image, namely depth map, mask image, and normal map. Furthermore, by adjusting the model to be cognizant of the subject’s geometry information, we can supervise NeRF generation in a more fine-grained manner, particularly by taking the shading mode into consideration.
While NeRF is renowned for effectively learning complex geometries, their proficiency in generating intricate texture details is hindered by significant memory demands. Therefore, our objective in the second stage is to enhance visual realism by transforming the coarse NeRF model into point clouds. Conversion to point clouds allows us to have the desirable groundtruth textures on the reference view through projection. Yet we observe that constructing point clouds from depth images as in [59] tends to introduce geometric inaccuracies since NeRF does not store any 3D geometry explicitly (only the density field). In lieu of this, our novel strategy entails an initial step of mesh generation, followed by the established mesh regularization techniques to refine the shape, before sampling points on the mesh surface. In addition, we harness the potential of the aforementioned subject-specific Text-to-Image (T2I) diffusion model, effectively elevating the texture realism, before extending the RGB rendering from the coarse NeRF to point clouds. This augmentation is shown to yield superior texture quality in the final outcomes.
Besides employing the proposed subject-specific 2D diffusion prior, we also utilize a 3D prior to further guide the training process as in [46], compensating the 3D information. Note that this 3D prior takes the reference image as input and thus both the adopted 2D and 3D prior are subject-specific knowledge priors to guide the novel views. To evaluate our method, we collect a benchmark consisting of 100 images. We compare Customize-It-3D with prior state-of-the-arts on RealFusion15 [38], and our Customize100 dataset. Extensive experiments demonstrate our method achieves significant improvement in visual quality in context of accurate geometry and realistic details. In summary, our contributions are:
  • We propose Customize-It-3D, an innovative framework for the generation of high-quality 3D content from a single image by utilizing a subject-specific diffusion prior, which is designed to enhance the personalization of 3D content creation for specific subjects.
  • We introduce a multi-modal DreamBooth model that promotes comprehensive knowledge priors derived from multi-modal images of the subject, prioritizing the faithfulness of the diffusion model to the reference image. This knowledge prior not only facilitates NeRF optimization based on the shading mode but also enhances the texture of the initial coarse results for better refinement.
  • The resulting framework demonstrates state-of-the-art performance in 3D reconstruction using a diverse range of in-the-wild images and images from existing datasets. Furthermore, this approach opens up intriguing possibilities, such as text-to-3D content creation.
2 Related work
Muti-view 3D reconstruction. Early works [1, 55, 14, 54] traditionally require a dense set of input images to accurately deduce geometry by establishing correspondence within the overlapping regions.. The advent of NeRF [39, 36] and its variants [41, 5, 6] have significantly propelled the quality of synthesized novel views through the reconstruction of 3D scenes via implicit neural representations. Subsequent efforts [20, 25, 11, 70] have been dedicated to the facilitation of NeRF optimization from a sparser set of input views. Typically they entail the extraction of per-view features from each input image and aggregate multi-view features for each point along the camera ray. This aggregated information is then decoded to determine density (or Signed Distance Function, SDF) and colors. Diffusion models, which have demonstrated significant advancements in general image synthesis, has emerged as a recent focal point in 3D generation research. Some studies endeavor to directly train 3D diffusion models based on diverse 3D representations, including point clouds [37, 42, 72, 75], meshes [15, 35], neural fields [2, 7, 9, 13, 17, 22, 24, 26, 28, 16, 43, 63, 73]. Others, such as [4, 65, 58, 57] conceptualize 3D generation as an image-to-image translation task and directly generate coherent multiview color images, as exemplified by Zero123 [34]. It’s noteworthy that both approaches require access to multi-view images for training. However, a recurring challenge in this domain lies in the constrained availability of expansive 3D assets datasets. Consequently, most studies have been confined to evaluation on limited categories of shapes or encounter difficulties in extending their models effectively on general objects.
3D generation using knowledge prior. A surge of approaches have been study leveraging large models pretrained on billions of images as knowledge prior to guide the 3D synthesis, circumventing the need for large-scale 3D datasets. One popular task is text-to-3D synthesis, given the exceptional breakthroughs in text-to-image generation. Pioneering works [40, 56] such as DreamFields [21], DreamFusion [45] and SJC [62] have demonstrated success of incorporating NeRF-like models with frozen generative systems for optimization-based 3D synthesis. They fed the rendered 2D images to diffusion models or CLIP model for calculating losses to guide the 3D shape optimization. It sparked numerous follow-up works improve such per-shape optimization scheme in the context of 3D representations [32, 60, 8, 30], sampling schedules [19], and loss design [64], subject-driven editing [48]. An alternative avenue of research [38, 59, 46, 33] seeks to generate 3D digital content from a single image, offering a more precise and controllable approach than text-based methods. The prevailing majority of existing endeavors in this area remain rooted in optimization-based techniques, achieving this task by imposing pixel-wise reconstruction losses on the reference view, in addition to inheriting the SDS loss from text-to-3D methodologies. In comparison, we also follow the DreamFusion pipeline [45], but with a focus on cultivating subject-specific knowledge prior. While Dreambooth3D [48] also leverages a subject-specific diffusion prior, it requires multiple subject images and is designed for text-to-3D editing. In contrast, our approach tackles a more challenging scenario with only one image and introduces a multi-modal diffusion prior, specifically aiming at image-to-3D generation.
Figure 2: We propose a two-stage framework for high-quality 3D creation from a reference image with subject-specific diffusion prior (Sec. 3.1). At the coarse stage, we optimize a NeRF for reconstructing the geometry of the reference image in a shading-aware manner(Sec. 3.2). We further build point clouds with enhanced texture from the coarse stage, and jointly optimize the texture of invisible points and a learnable deferred renderer to generate realistic and view-consistent textures (Sec. 3.3).
3 Method
Generating a personalized 3D object from a single image is a highly challenging task, because it involves disclosing the identity, geometric shape, and texture from limited and incomplete visual information. Therefore, we propose a two-stage coarse-to-fine framework, Customize-It-3D, leveraging carefully cultivated subject-specific knowledge prior to effectively constrain the coherency of 3D object with respect to a particular identity. The proposed framework is illustrated in Figure 2. We will first detail our subject-specific knowledge prior and then introduce our method.
3.1 Subject-Specific Knowledge Prior
Subject-Specific 2D Prior: Multi-modal DreamBooth. Current text-to-image diffusion models possess rich semantic knowledge and 2D image knowledge. DreamFusion [45] is a pioneering work that utilizes the latent knowledge of these diffusion models to guide a 3D representation optimization. However, due to the abundant imagination of 2D diffusion models, it is challenging to precisely control over the generated geometric shapes, textures, and identities using solely text, resulting in discrepancies and thereby causing confusion in guiding 3D reconstructions. This phenomenon tends to produce smoothed textures and imprecise geometries as a result of averaging the inconsistent outcomes. To address this issue, we undertake a fine-tuning process on a pre-trained text-to-image diffusion model using a single reference image. To mitigate the risks of overfitting and mode collapse, we choose to leverage multi-modal images in our approach.
Specifically, we first utilize a pre-trained monocular depth estimator [50] and a single-view normal estimator [12] [23] to obtain the depth and normal maps of the reference image. Furthermore, we segment the last channel of the input image to obtain a mask map. Then, we use this image set and the reference image together to partially fine-tune Stable Diffusion [51] with Dreambooth [52], so that it learns to bind a unique identifier with the specific subject embedded in the output domain of the model. We also find that a fully trained Dreambooth tends to overfit the subject viewpoints in input images, leading not only to a severe Janus problem but also reducing diversity, as noted in previous work [48]. Therefore, we only use the partially fine-tuned results. Unlike Dreambooth [52], we fine-tune not only the UNet but also the text encoder, which leads to better results.
Additionally, we utilize a prior preservation loss, which enables the supervision of model training during fine-tuning with self-generated images $\mathbf{x}_{ori}$ from the original diffusion model. Then our loss function to finetune a diffusion model $\bm{\epsilon}_{\phi}$ is :
$\begin{split}\mathbb{E}_{x,c,\epsilon,\epsilon^{\prime},t}[w_{t}||\bm{\epsilon}_{\phi}{(\alpha_{t}\mathbf{x}+\sigma_{t}\epsilon,c)-\mathbf{x}||_{2}^{2}}+\\ \lambda w_{t^{\prime}}||\bm{\epsilon}_{\phi}(\alpha_{t^{\prime}}\mathbf{x}_{ori}+\sigma_{t^{\prime}}\epsilon^{\prime},c_{ori})-\mathbf{x}_{ori}||_{2}^{2}],\end{split}$ (1)
where $\mathbf{x}$ is the reference image and $\alpha_{t},\sigma_{t},w_{t}$ control the noise schedule, $c_{ori}$ is the prompt to generate $x_{ori}$. The second term is prior-preservation loss with weight $\lambda$.
In terms of prompt processing, we need an identifier with a weak prior in both the language model and the diffusion model. Therefore, we follow previous work [61] and use the “sks” identifier for stable diffusion [51]. Differently, since we are dealing with a set of images with different modalities, it is imperative not to employ an identical prompt for all images, as this practice can introduce confusion to the model and lead it to erroneously emphasize common distinctive features. Our primary objective is not to model the shared aspects across all modalities but rather to equip the model with the capability to discern and accommodate the distinctions arising from different modalities. Thus, we incorporate additional instructions, and the prompt $c$ for fine-tuning the diffusion model takes the form of “a depth map / normal map / foreground mask / rgb photo of sks [class name]”. The choice of the specific instruction is made in accordance with the given modality during training. In this way, the model is encouraged to perceive the subject from different perspectives, enriching its subject-specific prior.
Subject-Specific 3D Prior: Viewpoint and Image-Conditioned Diffusion Model. Presently, large-scale pre-trained diffusion models are primarily trained on 2D images, lacking substantial 3D-specific knowledge. Notably, the recent Zero-1-to-3 model [34] stands out as it utilizes fine-tuning on a synthetic 3D dataset [10] to develop a viewpoint-conditioned diffusion model. This model takes a reference image along with related external camera parameters as inputs, enabling the generation of novel views of the same subject based on the reference image. Therefore, it serves as our preferred strategy for establishing a subject-specific 3D prior, contributing to improving 3D consistency.
3.2 Coarse Stage: Image to 3D Reconstruction
In the coarse stage, we aim to obtain the approximate geometric structure of the 3D object and the rough texture necessary for constructing a textured point cloud. In this stage, our scene model is a neural field based [39] on Instant-NGP [41]. This choice is made because it can handle complex topological changes in a smooth and continuous manner while also enabling the reconstruction of 3D objects in a relatively fast and computationally efficient manner.
Subject-specific knowledge prior for novel views. We optimize the neural radiance field such that its multi-view renderings look like high-quality samples using the subject-specific diffusion model. Specifically, let the rendered image $\mathbf{I}=\mathcal{G}_{\theta}(\mathbf{v})$ at the viewpoint $\mathbf{v}$, where $\mathcal{G}$ is the differentiable rendering function for NeRF optimization parameterized by $\theta$. We employ the multi-modal DreamBooth model as a 2D prior and follow [45]  [32] to use the score distillation (SDS) loss, which assigns a “score” to the rendered image, guiding the optimization of the 3D model’s parameters $\theta$ towards the direction of higher density regions.
Figure 3: The proposed shading mode-aware guidance largely enhances the 3D geometry in the coarse stage.
Differently, we go beyond this and propose modifying the text prompt based on the specific NeRF shading mode. This modification is aimed at fully harnessing the capabilities of our multi-modal diffusion model to offer precise guidance tailored to different NeRF shading modes, which is facilitated by the inherent multi-modal awareness of our finetuned diffusion model. To elaborate, given the text prompt $y$ that is generated from an image captioning model [31] from the reference image, we adapt the text prompt to “sks normal map of $y$” when the NeRF rendering employs “normal” shading mode and “sks rgb photo of $y$” when the NeRF rendering utilizes “albedo” shading mode. By enforcing consistency between the rendered image aligned with the modified text prompt, the multi-modal diffusion model is able to provide more accurate guidance and thus enhance the quality of the 3D reconstruction, as illustrated in Figure 3. Formally, we use 2D SDS loss as:
$\nabla_{\theta}\mathcal{L}_{SS2D}=\mathbb{E}_{t,\bm{\epsilon}}[w(t)(\bm{\epsilon}_{\phi}(\mathbf{z_{t}};y_{m},t)-\bm{\epsilon})\frac{\partial\mathbf{z}}{\partial\mathbf{I}_{m}}\frac{\partial\mathbf{I}_{m}}{\partial\theta}],$ (2)
where $\mathbf{I}_{m}=\mathcal{G}_{\theta}(\mathbf{v},m)$ and $m$ denotes the shading mode and can take on values of either “normal” or “albedo”. We use the personalized text prompt $y_{m}$ to encode NeRF’s novel view rendering $\mathbf{I}_{m}$ to the noisy latent $z_{t}$ by adding a random Gaussian noise $\bm{\epsilon}$ of a timestep $t$.
In addition, we have the subject-specific 3D prior for novel view guidance, which can be formulated as follows:
$\mathcal{L}_{SS3D}=\mathbb{E}_{t,\epsilon}\left[w(t)(\epsilon_{\phi}(\mathbf{z_{t};x},t,R,T)-\epsilon)\frac{\partial\mathbf{I}}{\partial\theta}\right],$ (3)
where $R,T$ represent the camera’s rotation and translation, and $\mathbf{x}$ stands for the input reference image. Here, we do not utilize text as guidance but instead use image-based guidance, specifically the relative camera transformations between viewpoints.
The overall supervision is:
$\nabla_{\theta}\mathcal{L}_{SS-SDS}=\lambda_{SS2D}\nabla_{\theta}\mathcal{L}_{SS2D}+\lambda_{SS3D}\nabla_{\theta}\mathcal{L}_{SS3D},$ (4)
where $\lambda_{SS3D}$ and $\lambda_{SS3D}$ are their weights respectively.
Groundtruth knowledge for the reference view. Under the reference view $\mathbf{v}_{ref}$ , the rendered image $\mathbf{I}=\mathcal{G}_{\theta}(\mathbf{v}_{ref})$ by NeRF should be consistent with the input image $\mathbf{x}$ to align with our goal of customizing 3D objects. Therefore, we utilize the pixel-wise difference between the NeRF rendering and the input image under the reference view as one of our major losses:
$\mathcal{L}_{ref}=||\mathbf{x}\odot M-\mathcal{G}_{\theta}(\mathbf{v}_{ref})||_{1},$ (5)
where $\odot$ is Hadamard product. And we follow previous work  [69] to apply foreground mask $M$ in order to get extracted object which will ease the geometry reconstruction.
Using only the RGB per-pixel losses can lead to poor quality geometry issues such as sunken faces, over-flattening, due to the inherent shape ambiguity in 3D reconstruction. To mitigate this issue, we employ MiDaS  [50], a pretrained monocular depth estimator, to estimate the depth $d_{ref}$ of the reference image. However, since the estimated pseudo depth may not be accurate and there is a scale and source mismatch with the depth $d$ from NeRF, we regularize the negative Pearson correlation between the pseudo depth and the depth from NeRF under the reference viewpoint,
$\mathcal{L}_{depth}=-\frac{\mathrm{Cov}(d_{ref},d)}{\mathrm{Var}(d_{ref})\mathrm{Var}(d)}.$ (6)
Here $\mathrm{Cov(.)}$ denotes covariance and $\mathrm{Var(.)}$ measures standard deviation. We employ this function as our depth regularization to encourage the NeRF output $d$ under the reference view to be close to the depth prior.
Figure 4: Illustrating different point cloud building methods from (1) depth images as in [59] and (2) our method.
Overall training. In summary, the loss in the coarse stage consists of the following parts: $\mathcal{L}_{ref}$, $\mathcal{L}_{SS-SDS}$ and $\mathcal{L}_{depth}$. During training, we follow [59], adopting a progressive training strategy. We start by training only in a partial region of the object’s front view, and then gradually expand the training scope.
3.3 Refine Stage: Neural Texture Enhancement
In the coarse stage, due to the computationally intensive nature of NeRF optimization, we only perform optimization at low resolutions. Moreover, due to the inherent limitations of NeRF, such as the tendency to produce high-frequency artifacts [45], we can only obtain a low-resolution, low-texture-quality 3D model. As a result, we are motivated to explore alternative representations for NeRF, particularly focusing on transforming it into an explicit point cloud. Point clouds have the advantage of allowing direct projection of images, preserving high-quality frontal texture details, and offering opportunities for personalization and enhancement of the projected images, fulfilling the needs of high-quality customized 3D models. In summary, our goal in the refine stage is to retain the geometric structure of the coarse stage, generate a dense point cloud, and optimize the textures not visible in the reference view $\mathbf{v}_{ref}$, as we directly project the frontal reference image $\mathbf{x}$ onto the point cloud. We elaborate several key designs to facilitate the transformation from the coarse to the refine stage.
Point cloud building. Past work has focused on constructing a point cloud from multi-view RGBD images under NeRF rendering [59]. However, due to noise in RGBD images and inaccurate depth estimation, the generated point cloud is very noisy and inaccurate, significantly impairing the 3D geometry in refine stage. To address this, we export the NeRF in coarse stage as a mesh, utilizing mesh regularizations and perform Poisson sampling on the mesh to obtain a dense point cloud $P^{m}$. Figure 4 shows this difference. We opt to a point cloud because it allows personalized diffusion model-based texture enhancement for each viewpoint, resulting in a higher-quality point cloud that better aligns with our customized 3D object generation needs.
Figure 5: The texture details from coarse rendering are largely enhanced using the multi-modal DreamBooth.
Texture projection. However, since NeRF-rendered images from different viewpoints may have overlapping regions, a 3D point may be assigned different colors under different viewpoints [66], leading to texture conflicts. Therefore, we propose an iterative strategy for constructing a clean texture point cloud $P$ from mesh sampled point cloud $P^{m}$. Initially, we trust all points $P^{m}_{ref}$ under reference view $\mathbf{v}_{ref}$ and use a reference mask $M$ to perform color mapping on $P^{m}_{ref}$, obtaining the front-facing perspective of the new point cloud $P_{ref}$. For novel view projection, we aim to avoid introducing points with color conflicts that overlap with $P_{ref}$. Hence, we proceed to project the recently obtained point cloud $P_{ref}$ onto each novel view $\mathbf{v}_{i}$ to generate a new mask denoted as $M^{ref\_proj}_{\mathbf{v}_{i}}$. We then project textures to the points under this novel view only when these points belong to the set difference between the novel view mask $M_{\mathbf{v}_{i}}$ and the reprojected mask $M^{ref\_proj}_{\mathbf{v}_{i}}$. Finally we have a clean point cloud $P=\{p_{ref},p_{1},...,p_{n}\}$ which ensures color mapping is done without introducing conflicts.
Figure 6: Illustrating the textured point cloud construction. We enhance the texture and mask of novel views and iteratively project texture to the point cloud.
Texture and mask enhancement. The texture quality obtained in the coarse stage is suboptimal. Therefore, prior to projecting the reference image $\mathbf{x}$ from the reference view and rendered RGB images $\mathcal{X}=\{\mathbf{x}_{\mathbf{v}_{1}},\mathbf{x}_{\mathbf{v}_{2}},...,\mathbf{x}_{\mathbf{v}_{n}}\}$ under novel views onto the point cloud geometry, we perform a texture enhancement process on all rendered images within $\mathcal{X}$. Specifically, we add noise to the rendered images and leverage the previously fine-tuned multi-modal personalized diffusion model to denoise into a set of pseudo images $\mathcal{X}^{pseudo}$ that possess higher texture quality as well as multi-view consistency. Figure 5 shows that the texture details can be significantly improved by utilizing our multi-modal DreamBooth model.
Simultaneously, due to the issues in the coarse stage of NeRF, the generated masks =$\mathcal{M}=\{M_{\mathbf{v}_{1}},M_{\mathbf{v}_{2}},...,M_{\mathbf{v}_{n}}\}$ by NeRF are not satisfactory. Fine details (for example long and thin parts) in the object structure may be compromised, leading to defects in point cloud texture and geometry initialization due to mask inadequacies. Therefore, we adopt Segment Anything Model (SAM) [27] to generate more refined and accurate masks $\mathcal{M}^{pseudo}$. However, due to the insufficient resolution of RGB images and the noise introduced by SAM, we assess the quality of the generated mask by relying on SAM’s score. Only when the quality surpasses a certain threshold do we consider using this mask; otherwise, we adhere to utilizing masks generated by NeRF. We illustrate our textured point cloud building in Figure 6.
Figure 7: Qualitative comparison on image-to-3D generation. We compare Customize-It-3D to RealFusion [38], Make-it-3D [59] and Magic123 [46] for creating 3D objects from a single unposed image (the leftmost column).
Deferred point cloud rendering and training. After initializing the geometry and texture of the point cloud $P$, a focus is placed on enhancing the texture of the point cloud in novel views with subject-specific prior. Specifically, we assign a 19-dimensional descriptor $F$ to each point in the novel view $\mathbf{v}$, where the first three dimensions are colors initialized from the above process. During the rendering of the point cloud, we employ a deferred rendering scheme [59], rendering the point cloud K times at different resolutions and concatenating the resulting K feature maps into the final image using a jointly optimized U-Net network denoted as $\mathcal{R}_{\theta}$. The refine stage’s loss is similar to that in Sec. 3.2, with the addition of a regularization term penalizing excessive differences in texture between the refined and initialized states.
4 Experiments
4.1 Implementation Details
[…]
Table 1: Quantitative comparison on RealFusion15. We compute LPIPS and PSNR under the reference view, and CLIP under novel views.
[…]
NeRF and point cloud rendering. We employ multi-scale hash encoding from Instant-NGP [41] and maintain an occupancy grid to skip empty space for improving rendering efficiency. Additionally, we incorporate shading augmentations, such as Lambertian shading similar to [45]. However, during the first 1000 iterations of the second phase of progressive training, we utilize normal shading to emphasize the learning of better geometry. For the deferred rendering strategy of point clouds, we utilize a 2D U-Net architecture with gated convolutions [71].
4.2 Comparisons with the State of the Arts
Baselines. We compare our approach against recent state-of-the-art methods: RealFusion [38], Make-It-3D [59], and Magic123  [46] (using Zero123 XL checkpoint  [34] as ours for fair comparison). We compare these methods using the Realfusion dataset [38] and a customized dataset we created. The Realfusion dataset includes many natural images, while our customized dataset comprises real images and images generated by Stable Diffusion XL [44]. We evaluate all the baseline methods using their official code.
Figure 8: The effect of Multi-modal DreamBooth, showing a consistent improvement in terms of both geometry and texture.
Figure 9: Single RGB versus Multi-modal DreamBooth, showing a clear improvement particularly in geometry.
Qualitative comparison. We present a comprehensive collection of qualitative results in Figure 7. RealFusion often generates flat 3D results with colors and shapes that diverge significantly from the input image. Make-it-3D exhibits competitive texture quality but suffers from an issue known as long geometry in side views, particularly noticeable in the reconstruction of objects such as chairs. Magic123 produces visually plausible structures but grapples with a notable issue of multi-face, as it tends to replicate the reference texture in the back view.. In contrast, our approach can reasonably hallucinate the texture details and geometry for novel views that even deviate significantly from the reference image, which greatly improves the fidelity and consistency in creating 3D models.
Quantitative comparison. We quantitatively compare with the baselines in Table 1. We use metrics following [59, 46] with PSNR, LPIPS [74], and CLIP-similarity [47]. As shown in the table, Customize-It-3D achieves Top-1 performance across all the metrics, showing that our model is able to generate 3D objects with better 3D consistency.
5 Ablations and Analysis
With or without multi-modal DreamBooth. We first ablate the effect of using the proposed multi-modal Dreambooth in Figure 8. We observe a consistent improvement in terms of both texture and geometry. Employing the suggested multi-modal subject-specific diffusion model prior results in higher-quality 3D content, yielding more compelling visuals with enhanced 3D consistency.
Single RGB versus multi-modal. We experiment a baseline that employs conventional DreamBooth finetuned from a single RGB reference to further investigate the impact of our shading mode-aware multi-modal guidance. The difference shown in Figure 9 demonstrate a clear improvement, whereas the baseline suffers from multi-face problem.
Figure 10: Illustrating high-fidelity text-to-3D generation.
Figure 11: Illustrating high-fidelity 3D creation for real scenes.
6 Applications
High-quality text-to-3D generation. To achieve high-quality text-to-3D creation, we initially utilize a T2I diffusion model to convert the text prompt into a reference image and then pass it into our image-based 3D creation method. We show the results in Figure 10. Customize-It-3D demonstrates remarkable quality in text-to-3D generation.
Real scene modeling. We extend our model evaluation to real-world scenarios by inputting arbitrary general images capturing complex, authentic scenes into our framework. Figure 11 presents several visual examples, verifying the efficacy of our method in modeling real scenes with a high degree of capability and fidelity.
7 Conclusion
We have presented Customize-It-3D, a novel two-stage approach for image-to-3D generation. Leveraging the established customized knowledge prior, we enhances NeRF optimization for improved geometry and refines texture, resulting in superior alignment of 3D content with the subject. Customize-It-3D demonstrates versatility in handling general objects and empowers captivating applications.
Supplementary Material
Figure 12: Additional comparison with baseline: RealFusion [38], Make-It-3D [59], and Magic123  [46]. The first column is the reference image.
A Additional Implementation Details
A.1 Coarse stage
Multi-modal Dreambooth finetuning and usage. We utilized the stable diffusion V2.0 model [51] in our experiments. We employed multi-modal images related to the reference image and conducted a four-step fine-tuning process on the diffusion model. The sequence and number of iterations for each training step are as follows: 25 iterations for mask image training, 50 iterations for depth image training, 50 iterations for normal image training, and 75 iterations for RGB image training, totaling 200 iterations. During fine-tuning, prompts were specified as follows: “a foreground mask / depth map / normal map / RGB photo of sks [class_name].” When using the finetuned model for training in the coarse stage, we modified the prompts based on shading mode. For normal shading, we append the modifier “normal map” to the text prompt, while for other shading strategies, we add the modifier “RGB photo”. We observed that this approach facilitated the better utilization of subject-specific knowledge, leading to improved results.
Optimizing the pipeline. We employ the Adam optimizer [67] with a learning rate of 0.005 and no weight decay for 7000 iterations. Simultaneously, during the coarse stage, we implement progressive training, where the first 2000 iterations exclusively trained within the range of ±$\pm 45^{\circ}$ from the reference view, and the subsequent 5000 iterations covered the entire $360^{\circ}$ around the object. In the refine stage, we conducted 3000 iterations to optimize the texture.
Camera setting. For the camera sampling method, we randomly sample novel views with a probability of 75%, while with a 25% probability, we sample a reference view following the approach in [45]. The reference view is assumed to be a frontal view, i.e., a polar angle of $90^{\circ}$ and an azimuth angle of $0^{\circ}$. It is further assumed that the camera is positioned at a distance of 1 meter from the origin of the coordinate system, and the field of view (FOV) is set to $60^{\circ}$.
A.2 Refine stage
Post-mask processing. We also perform post-mask-processing, wherein we apply certain morphological operations to both SAM-generated and NeRF-generated masks, i.e., taking the union of the two masks to complement fine structures lacking in the initial mask. Finally, we have $M_{pseudo}=\hat{M}\cup\hat{M}_{SAM}$, where $\hat{M}$ and $\hat{M}_{sam}$ are the rendering and SAM-generated mask after post-processing under novel view.
Point cloud building. The goal of our point cloud building strategy is to avoid introducing points with color conflicts that overlap with existing point cloud $p^{m}_{ref}$, which is build from reference view. We use a reference mask $M^{ref}_{pseudo}$ to perform color mapping on $p^{m}_{ref}$, obtaining the front-facing perspective of the new point cloud $p_{ref}$. For all remaining regions under novel view, we project the just obtained point cloud $p_{ref}$ onto each novel view $v_{i}$ to create a new mask $m^{re}_{i}$. We then take the set difference between the novel view mask $m^{i}_{pseudo}$ and the re-projected mask $m^{re}_{i}$ as the mask for the remaining points $P_{left}=P^{m}-p^{m}_{ref}$. Finally we will have a clean point cloud $P=\{p_{ref},p_{1},...,p_{n}\}$ which ensures color mapping is done without introducing conflicts.
Point cloud rendering. In order to perform better point cloud rendering, each point’s has a 19-dimensional descriptor, with the initial three dimensions initialized with RGB colors and the rest initialized randomly [59]. Deferred rendering is conducted for a total of $K=3$ times, with resolutions denoted as $[H/2^{i},W/2^{i}]$, where $i\in[0,K)$ and ultimately, the learned U-Net network concatenates these three feature maps. And the U-Net architecture contains 3 down-sampling and up-sampling layers with gated convolutions[71]. The rendering resolution is set to $128\times 128$ in the coarse stage for NeRF and $800\times 800$ in the refine stage for the point cloud, respectively.
Figure 13: Additional results of an ablation study on whether to use the multi-modal dreambooth. The first column is the reference image.
Figure 14: Additional ablation results. The first column is the reference image. The second column displays the results obtained using our method. The third column showcases the outcomes when substituting the multi-modal dreambooth with the traditional dreambooth. The fourth column exhibits the results when shading mode guidance is omitted during the NeRF training process.
B Additional Ablation Study Results
As mentioned in the paper, current methods for text/image-to-3D rely on a generic diffusion model with strong generalization capabilities, leveraging implicit 2D knowledge to guide the generation of novel views. However, due to its overly rich imagination, this model may generate 3D objects that do not align with the subject identity. To address this limitation, we fine-tune the diffusion model to learn subject-oriented knowledge. We conduct a series of experiments to demonstrate the effectiveness of our subject-specific and multi-modal diffusion model. We first ablate the effect of using the proposed multi-modal Dreambooth, as illustrated in Figure 13. The results indicate that without the guidance of our multi-modal Dreambooth, the coarse stage produces textures and geometry that deviate from the subject identity, thereby impacting the initialization of point cloud geometry and textures in the refine stage, as exemplified by the patterns on the fox’s neck and the rabbit’s ears.
Furthermore, we compare the results obtained using the traditional dreambooth (trained with a single RGB image) with our multi-modal dreambooth (trained with multiple multi-modal images). We also compare the results of the NeRF training process with and without shading mode guidance. Both sets of experiments are depicted in Figure 14. The results show that a simple attempt using only dreambooth leads to overfitting and mode collapse. Additionally, we observed that using shading mode guidance better leverages our multi-modal dreambooth, resulting in more convergent and 3D-consistent outcomes.
[…]
Table 2: Quantitative comparison on our dataset. We compute LPIPS, PSNR and CLIP-similarity.
C Additional Comparisons with Baseline
In this section, we show more results comparing our method with the baseline: RealFusion [38], Make-It-3D [59], and Magic123  [46] (using Zero123 XL checkpoint [34]), as shown in the Figure 12. Also we do quantitative comparison on our own dataset as shown in the Table 2, and our method has the best results. Results show that our method exhibits increased fidelity to the input reference image in both geometry and texture, thanks to the utilization of subject-specific knowledge priors. This incorporation makes the training process more aligned with the subject’s identity. Particularly in terms of texture, as exemplified in the second row with the silver cat in the Figure 12, our method generates back textures with a more pronounced metallic sheen, consistent with the characteristics of the front texture. Similarly, as seen in the fifth row with the doll in the Figure 12, our method generates doll textures on the back and sides, including patterns on the skirt and hair color, that better align with the subject’s identity. Therefore, our method is capable of generating textures that are more consistent with the reference view.
D Additional Results
In this section, we show more results using our method. Figure 15 and Figure 16 demonstrates the ability of our method to generate high-quality 3D objects from a single image.
E Limitation
Our method has several limitations. Firstly, because our approach relies on using pretrained models as depth estimator [50] and normal estimator [12] [23] to estimate the depth map and normal map of the reference image. So any errors in these modules will affect our method and impact the overall generation quality. Also, our method, relying on generative prior to imagine the 3D content from the single reference viewpoint, inevitably introduces geometry ambiguity [45], leading to issues such as the Janus problem or over-flat geometry.
Figure 15: Additional results by Customize-It-3D. The first column is the reference image.
Figure 16: Additional results by Customize-It-3D. The first column is the reference image.