### v3

# High-Quality 3D Creation from A Single Image Using Subject-Specific Knowledge Prior

Abstract
In this paper, we address the critical bottleneck in robotics caused by the scarcity of diverse 3D data by presenting a novel two-stage approach for generating high-quality 3D models from a single image. This method is motivated by the need to efficiently expand 3D asset creation, particularly for robotics datasets, where the variety of object types is currently limited compared to general image datasets. Unlike previous methods that primarily rely on general diffusion priors, which often struggle to align with the reference image, our approach leverages subject-specific prior knowledge. By incorporating subject-specific priors in both geometry and texture, we ensure precise alignment between the generated 3D content and the reference object. Specifically, we introduce a shading mode-aware prior into the NeRF optimization process, enhancing the geometry and refining texture in the coarse outputs to achieve superior quality. Extensive experiments demonstrate that our method significantly outperforms prior approaches.
I INTRODUCTION
The development of robotic models is increasingly constrained by a critical bottleneck: the limited diversity of 3D data available for training. While general image datasets, such as ImageNet[10], contain thousands of object categories, robotics datasets are far less diverse. For instance, ShapeNet[5] and GraspNet[14], two commonly used datasets, include only 55 and 88 object categories. This limitation restricts the ability of robots to effectively interact with a wide range of real-world objects. To address this, there is an urgent need to expand the variety of 3D assets in robotics datasets, enabling more robust and versatile policy models.
At the same time, we also aim to leverage image-to-3D techniques to improve robot interactions with real-world objects. For example, when a robot captures an image of an object using its 2D camera, we can convert that image into a 3D model, thereby enhancing the robot’s ability to interact with and manipulate the object. This approach has the potential to significantly improve robotic performance in complex, dynamic environments.
However, this task remains deeply intricate due to its inherently ill-posed nature, as there is a substantial domain gap between a single 2D image and the full 3D spatial and textural characteristics of the object. Despite its great challenge, recent works [64, 49, 40, 72] have achieved notable results in image-to-3D creation. They draw inspiration from text-to-3D methods [48, 68] which utilize score distillation sampling (SDS) loss to optimize a neural radiance field (NeRF). While producing promising results, they still exhibit noticeable inconsistencies when compared with the reference view, either in context of the geometry (i.e., multi-face issue) or from the texture perspective (novel views often lack fidelity, presenting smooth and disruptive particulars).
Despite incorporating several adjustments such as CLIP loss and textual inversion considered in [64, 49], the supervision of novel views still relies on a general text-to-image model [52, 54, 56, 3, 19], which has challenge in producing consistent results based solely on text, given the inherent difficulty in crafting textual descriptions to cover every detail. Therefore, these approaches struggle to maintain consistency and fall short in reconstructing high-fidelity 3D objects. This realization has inspired us to explore strategies for cultivating subject-specific knowledge prior, aiming to achieve high-quality image-to-3D.
We employ a two-stage framework, which is a widely recognized way for achieving high-quality image-to-3D. We introduce a subject-specific diffusion prior, customized by the reference image, to fully utilize the geometric and appearance information embedded within the reference image. Building on this prior, we propose shading mode-aware guidance in the NeRF optimization process, and texture enhancement in the refinement stage. We name our approach Customize-It-3D, reflecting the utilization of the proposed customized prior.
In the first stage, we optimize a neural radiance field to learn an implicit volume representation. The reference view is directly supervised through rendering loss, specifically pixel differences. For novel views, unlike conventional approaches employing general T2I models for SDS loss, we use a subject-specific T2I model. There has been remarkable success in subject-driven image generation [55, 30]. They typically demand at least 3 to 5 images capturing the same subject with varying contexts in order to model the key visual features of the subject. However, our scenario presents a challenge as we possess only a solitary image, which may misguide the model to a trivial solution (mode collapse). To surmount this limitation, we propose to exploit multiple modalities extracted from the reference image, namely depth map, mask image, and normal map. More importantly, by adjusting the model to incorporate the subject’s geometric information, we can more precisely guide the generation of NeRF, especially by considering the shading mode in our supervision.
While NeRF is renowned for effectively learning complex geometries, their proficiency in generating intricate texture details is hindered by significant memory demands. Therefore, our objective in the second stage is to enhance visual realism by transforming the coarse NeRF model into point clouds. Conversion to point clouds allows us to have the desirable groundtruth textures on the reference view through projection. Yet we observe that constructing point clouds from depth images as in [64] tends to introduce geometric inaccuracies since NeRF does not store any 3D geometry explicitly (only the density field). In light of this, our strategy entails an initial step of mesh generation, followed by the established mesh regularization techniques to refine the shape, before sampling points on the mesh surface. In addition, we harness the potential of the aforementioned subject-specific diffusion model, to effectively elevate the texture realism, before extending the RGB rendering from the coarse NeRF to point clouds. This augmentation is shown to yield superior texture quality in the final outcomes.
To evaluate our method, we collect a benchmark including 100 images, and compare with baselines on RealFusion [40], and our dataset. In summary, our contributions are:
  • We propose Customize-It-3D, an innovative framework designed to address the critical bottleneck in robotics caused by the limited diversity of 3D data. By generating high-quality 3D content from a single image using a subject-specific diffusion prior, this framework enhances the personalization of 3D asset creation, which is essential for expanding object variety in robotics datasets.
  • We introduce a multi-modal DreamBooth model that promotes comprehensive knowledge priors derived from multi-modal images of the subject, prioritizing the faithfulness of the diffusion model to the reference image. This knowledge prior not only facilitates shading mode-aware NeRF optimization but also enhances the texture for better refinement.
  • The resulting framework demonstrates state-of-the-art performance in 3D reconstruction using a diverse range of in-the-wild images and images from existing datasets.
II Related work
Muti-view 3D reconstruction. Early works [1, 58, 15, 57] traditionally require multiple input images to deduce geometry by establishing correspondence within the overlapping regions. The advent of NeRF [41, 37, 43, 6, 7] have significantly propelled the quality via implicit neural representations. Subsequent efforts [22, 26, 11, 75] have been dedicated to facilitate NeRF optimization from a sparser set of input views. Typically they entail the extraction of per-view features from each input image and aggregate multi-view features for each point. This aggregated information is then decoded to determine density (or Signed Distance) and colors. Diffusion models have emerged as a recent focal point in 3D generation research. Some studies endeavor to directly train 3D diffusion models based on diverse 3D representations, including point clouds [39, 44, 76, 79], meshes [16, 36], neural fields [2, 8, 9, 13, 18, 23, 25, 27, 29, 17, 45, 67, 77]. Others, such as [4, 69, 62, 60] conceptualize 3D generation as an image-to-image translation and directly generate coherent multiview images, as exemplified by Zero123 [34].
3D generation using knowledge prior. Pioneering works [42, 59] such as DreamFields [21], DreamFusion [48] and SJC [65] sparked numerous follow-up works which [40, 64, 49, 33, 66, 70, 38, 33, 61] seeks to generate 3D digital content from a single image, offering a more precise and controllable approach. There are two categories along this direction. One is data-driven approach that trains a feed-forward model with extensible 3D supervision. For instance, LRM [20] adopts a large transformer-based encoder-decoder architecture for learning 3D representations of objects from a single image. DMV3D [73] trains on large scale multi-view image datasets of highly diverse object. Another existing endeavors in this area remain rooted in optimization-based techniques, imposing pixel-wise reconstruction losses on the reference view, in addition to the SDS loss. For example, SyncDreamer [35] adopts a synchronized multiview diffusion to model the joint probability distribution of multiview images for multiview consistency. Magic123 [49] uses 2D and 3D priors simultaneously to generate faithful 3D content from any given image. DreamGaussian [63] designs a generative 3D Gaussian Splatting model with companioned mesh extraction and texture refinement in UV space, aiming at accelerating the optimization process.
In comparison, we follow the optimization-based pipeline [48], but focus on cultivating subject-specific knowledge prior. While Dreambooth3D [51] also leverages customized diffusion prior, it requires multiple subject images and is designed for text-to-3D editing. Instead, our approach tackles a more challenging scenario with only one image and introduces a multi-modal diffusion prior.
Figure 1: We propose a two-stage framework for 3D creation from a reference image with subject-specific diffusion prior (Sec. III-A). At the coarse stage, we optimize a NeRF for reconstructing the geometry of the reference image in a shading-aware manner(Sec. III-B). We further build point clouds with enhanced texture from the coarse stage, and jointly optimize the texture of invisible points and a learnable deferred renderer to generate realistic and view-consistent textures (Sec. III-C).
III Method
In this work, we propose to leverage carefully cultivated subject-specific knowledge prior to effectively constrain the coherency of 3D object to a particular identity to address the need for expanding 3D assets in robotics. The proposed framework is illustrated in Figure 1.
III-A Subject-Specific Knowledge Prior
Subject-Specific Prior: Multi-modal DreamBooth. Current text-to-image diffusion models possess rich semantic knowledge and 2D image knowledge. DreamFusion [48] is a pioneering work that utilizes the latent knowledge of these diffusion models to guide a 3D representation optimization. However, due to the abundant imagination of 2D diffusion models, it is challenging to precisely control over the generated geometric shapes, textures, and identities using solely text, resulting in discrepancies and thereby causing confusion in guiding 3D reconstructions. This phenomenon tends to produce smoothed textures and imprecise geometries as a result of averaging the inconsistent outcomes. To address this issue, we undertake a fine-tuning process on a pre-trained text-to-image diffusion model using a single reference image. To mitigate the risks of overfitting and mode collapse, we choose to leverage multi-modal images in our approach.
Specifically, we first utilize a pre-trained monocular depth estimator [53] and a single-view normal estimator [12] [24] to obtain the depth and normal maps of the reference image. Furthermore, we segment the last channel of the input image to obtain a mask map. Then, we use this image set and the reference image together to partially fine-tune Stable Diffusion [54] with DreamBooth [55], so that it learns to bind a unique identifier with the specific subject embedded in the output domain of the model. We also find that a fully trained DreamBooth tends to overfit the subject viewpoints in input images, leading not only to a severe Janus problem but also reducing diversity, as noted in previous work [51]. Therefore, we only use the partially fine-tuned results. Unlike DreamBooth [55], we fine-tune not only the UNet but also the text encoder, which leads to better results.
Additionally, we utilize a prior preservation loss, which enables the supervision of model training during fine-tuning with self-generated images $\mathbf{x}_{o}$ from the original diffusion model. The loss is,
$\begin{split}\mathbb{E}_{x,c,\epsilon,\epsilon^{\prime},t}[w_{t}||\bm{\epsilon}_{\phi}{(\alpha_{t}\mathbf{x}+\sigma_{t}\epsilon,c)-\mathbf{x}||_{2}^{2}}+\\ \lambda w_{t^{\prime}}||\bm{\epsilon}_{\phi}(\alpha_{t^{\prime}}\mathbf{x}_{o}+\sigma_{t^{\prime}}\epsilon^{\prime},c_{o})-\mathbf{x}_{o}||_{2}^{2}],\end{split}$ (1)
where $\mathbf{x}$ is the reference image and $\alpha_{t},\sigma_{t},w_{t}$ control the noise schedule, $c_{o}$ is the prompt to generate $x_{ori}$ and c is the prompt with a unique identifier for reference image.
In terms of prompt processing, we need an identifier with a weak prior in both the language model and the diffusion model. Therefore, we follow previous work [46] and use the “sks” identifier for stable diffusion [54]. Differently, since we are dealing with a set of images with different modalities, it is imperative not to employ an identical prompt for all images, as this practice can introduce confusion to the model and lead it to erroneously emphasize common distinctive features. Our primary objective is not to model the shared aspects across all modalities but rather to equip the model with the capability to discern and accommodate the distinctions arising from different modalities. Thus, we incorporate additional instructions, and the prompt $c$ for fine-tuning the diffusion model takes the form of “a depth map / normal map / foreground mask / rgb photo of sks [class name]”. The choice of the specific instruction is made in accordance with the given modality during training. In this way, the model is encouraged to perceive the subject from different perspectives.
III-B Coarse Stage: Image to 3D Reconstruction
In the coarse stage, we aim to obtain the approximate geometric structure of the 3D object and the rough texture necessary for constructing a textured point cloud. In this stage, our scene model is a neural field based [41] on Instant-NGP [43]. This choice is made because it can handle complex topological changes in a smooth and continuous manner while also enabling the reconstruction of 3D objects in a relatively fast and computationally efficient manner.
Figure 2: An example of a Lego castle. Point cloud building results from (1) depth images [64] and (2) our method.
Subject-specific knowledge prior for novel views. We optimize the neural radiance field using the subject-specific diffusion model. Specifically, let the rendered image $\mathbf{I}=\mathcal{G}_{\theta}(\mathbf{v})$ at the viewpoint $\mathbf{v}$, where $\mathcal{G}$ is the differentiable rendering function for NeRF optimization parameterized by $\theta$. We employ the multi-modal DreamBooth model as a 2D prior and follow [48]  [32] to use the score distillation (SDS) loss, which assigns a “score” to the rendered image, guiding the optimization of the 3D model’s parameters $\theta$ towards the direction of higher density regions.
Differently, we go beyond this and propose modifying the text prompt based on the specific NeRF shading mode. This modification is aimed at fully harnessing the capabilities of our multi-modal diffusion model to offer precise guidance tailored to different NeRF shading modes, which is facilitated by the inherent multi-modal awareness of our finetuned diffusion model. To elaborate, given the text prompt $y$ that is generated from an image captioning model [31] from the reference image, we adapt the text prompt to “sks normal map of $y$" when the NeRF rendering employs “normal" shading mode and “sks rgb photo of $y$" when the NeRF rendering utilizes “albedo" shading mode. By enforcing consistency between the rendered image aligned with the modified text prompt, the multi-modal diffusion model is able to provide more accurate guidance and thus enhance the quality of the 3D reconstruction. Formally, we use 2D SDS loss as:
$\nabla_{\theta}\mathcal{L}_{SS2D}=\mathbb{E}_{t,\bm{\epsilon}}[w(t)(\bm{\epsilon}_{\phi}(\mathbf{z_{t}};y_{m},t)-\bm{\epsilon})\frac{\partial\mathbf{z}}{\partial\mathbf{I}_{m}}\frac{\partial\mathbf{I}_{m}}{\partial\theta}],$ (2)
where $\mathbf{I}_{m}=\mathcal{G}_{\theta}(\mathbf{v},m)$ and $m$ denotes the shading mode and can take on values of either “normal" or “albedo". We use the personalized text prompt $y_{m}$ to encode NeRF’s novel view rendering $\mathbf{I}_{m}$ to the noisy latent $z_{t}$ by adding a random Gaussian noise $\bm{\epsilon}$ of a timestep $t$.
In addition, we leverage a 3D prior for novel view guidance, thanks to the recent Zero-1-to-3 model [34]. The 3D loss can be formulated as follows:
$\nabla_{\theta}\mathcal{L}_{SS3D}=\mathbb{E}_{t,\epsilon}\left[w(t)(\epsilon_{\phi}(\mathbf{z_{t};x},t,R,T)-\epsilon)\frac{\partial\mathbf{I}}{\partial\theta}\right],$ (3)
where $R,T$ represent the camera’s rotation and translation, and $\mathbf{x}$ stands for the input reference image. The overall supervision is:
$\nabla_{\theta}\mathcal{L}_{SS-SDS}=\lambda_{SS2D}\nabla_{\theta}\mathcal{L}_{SS2D}+\lambda_{SS3D}\nabla_{\theta}\mathcal{L}_{SS3D},$ (4)
where $\lambda_{SS3D}$ and $\lambda_{SS3D}$ are their weights respectively.
Groundtruth knowledge for the reference view. Under the reference view $\mathbf{v}_{ref}$ , the rendered image $\mathbf{I}=\mathcal{G}_{\theta}(\mathbf{v}_{ref})$ by NeRF should be consistent with the input image $\mathbf{x}$ to align with our goal of customizing 3D objects. Therefore, we utilize the pixel-wise difference between the NeRF rendering and the input image under the reference view as one of our major losses:
$\mathcal{L}_{ref}=||\mathbf{x}\odot M-\mathcal{G}_{\theta}(\mathbf{v}_{ref})||_{1},$ (5)
where $\odot$ is Hadamard product. And we follow previous work  [74] to apply foreground mask $M$ in order to get extracted object which will ease the geometry reconstruction.
Using only the RGB per-pixel losses can lead to poor quality geometry issues such as sunken faces, over-flattening, due to the inherent shape ambiguity in 3D reconstruction. To mitigate this issue, we employ MiDaS  [53], a pretrained monocular depth estimator, to estimate the depth $d_{ref}$ of the reference image. However, since the estimated pseudo depth may not be accurate and there is a scale and source mismatch with the depth $d$ from NeRF, we regularize the negative Pearson correlation between the pseudo depth and the depth from NeRF under the reference viewpoint,
$\mathcal{L}_{depth}=-\frac{\mathrm{Cov}(d_{ref},d)}{\mathrm{Var}(d_{ref})\mathrm{Var}(d)}.$ (6)
Here $\mathrm{Cov(.)}$ denotes covariance and $\mathrm{Var(.)}$ measures standard deviation. We employ this function as our depth regularization to make the NeRF output $d$ under the reference view close to the depth prior.
Overall training. In summary, the loss in the coarse stage consists of $\mathcal{L}_{ref}$, $\mathcal{L}_{SS-SDS}$ and $\mathcal{L}_{depth}$. During training, we follow [64], adopting a progressive training strategy. We start by training only in a partial region of the object’s front view, and then gradually expand the training scope.
Figure 3: Qualitative comparison on image-to-3D generation. We compare Customize-It-3D to RealFusion [40], Make-it-3D [64] , Magic123 [49] and DreamGaussian [63] for creating 3D objects from a single unposed image (the leftmost column).
Figure 4: Comparison with Magic123 in coarse stage.
III-C Refine Stage: Neural Texture Enhancement
In the coarse stage, due to the computationally intensive nature of NeRF optimization, we only perform optimization at low resolutions. Moreover, due to the inherent limitations of NeRF, such as the tendency to produce high-frequency artifacts [48], we can only obtain a low-resolution, low-texture-quality 3D model. As a result, we are motivated to explore alternative representations for NeRF, particularly focusing on transforming it into an explicit point cloud. Point clouds have the advantage of allowing direct projection of images, preserving high-quality frontal texture details, and offering opportunities for personalization and enhancement of the projected images, fulfilling the needs of high-quality customized 3D models. In summary, our goal in the refine stage is to retain the geometric structure of the coarse stage, generate a dense point cloud, and optimize the textures not visible in the reference view $\mathbf{v}_{ref}$, as we directly project the frontal reference image $\mathbf{x}$ onto the point cloud. We elaborate several key designs to facilitate this transformation.
Point cloud building. Past work has focused on constructing a point cloud from multi-view RGBD images under NeRF rendering [64]. However, due to noise in RGBD images and inaccurate depth estimation, the generated point cloud is very noisy , significantly impairing the 3D geometry in refine stage. To address this, we export the NeRF in coarse stage as a mesh, utilizing mesh regularizations and perform Poisson sampling on the mesh to obtain a dense point cloud $P^{m}$. Figure 2 shows this difference. We opt to a point cloud because it allows personalized diffusion model-based texture enhancement for each viewpoint, resulting in a higher-quality point cloud that better aligns with our customized 3D object generation needs.
Texture projection. However, since NeRF-rendered images from different viewpoints may have overlapping regions, a 3D point may be assigned different colors under different viewpoints [71], leading to texture conflicts. Therefore, we propose an iterative strategy for constructing texture point cloud $P$ from mesh sampled point cloud $P^{m}$. Initially, we trust all points $P^{m}_{ref}$ under reference view $\mathbf{v}_{ref}$ and perform color mapping on $P^{m}_{ref}$, obtaining the front-facing perspective of the new point cloud $P_{ref}$. For novel view projection, we aim to avoid introducing points with color conflicts that overlap with $P_{ref}$. Hence, we proceed to reproject the recently obtained point cloud $P_{ref}$ onto each novel view $\mathbf{v}_{i}$ to generate a new mask denoted as $M^{ref\_proj}_{\mathbf{v}_{i}}$. We then project textures to the points under this novel view only when these points belong to the set difference between the novel view mask $M_{\mathbf{v}_{i}}$ and the reprojected mask $M^{ref\_proj}_{\mathbf{v}_{i}}$. Finally we have a clean point cloud $P=\{p_{ref},p_{1},...,p_{n}\}$ which ensures color mapping is done without introducing conflicts.
Texture and mask enhancement. The texture quality obtained in the coarse stage is suboptimal. Therefore, prior to projecting the reference image $\mathbf{x}$ from the reference view and rendered RGB images $\mathcal{X}=\{\mathbf{x}_{\mathbf{v}_{1}},\mathbf{x}_{\mathbf{v}_{2}},...,\mathbf{x}_{\mathbf{v}_{n}}\}$ under novel views onto the point cloud geometry, we perform a texture enhancement process on all rendered images within $\mathcal{X}$. Specifically, we add noise to the rendered images and leverage the previously fine-tuned multi-modal personalized diffusion model to denoise into a set of pseudo images $\mathcal{X}^{pseudo}$ that possess higher texture quality as well as multi-view consistency. Simultaneously, due to the issues in the coarse stage of NeRF, the generated masks =$\mathcal{M}=\{M_{\mathbf{v}_{1}},M_{\mathbf{v}_{2}},...,M_{\mathbf{v}_{n}}\}$ by NeRF are not satisfactory. Fine details (for example long and thin parts) in the object structure may be compromised, leading to defects in point cloud texture and geometry initialization due to mask inadequacies. Therefore, we adopt Segment Anything Model (SAM) [28] to generate more refined and accurate masks $\mathcal{M}^{pseudo}$. However, due to the insufficient resolution of RGB images and the noise introduced by SAM, we assess the quality of the generated mask by relying on SAM’s score. Only when the quality surpasses a certain threshold do we consider using this mask; otherwise, we adhere to utilizing masks generated by NeRF.
Figure 5: The effect of multi-modal DreamBooth.
Figure 6: Our proposed method multi-modal DreamBooth with shading mode-aware guidance shows a clear improvement. All results are obtained under novel view in the coarse stage.
(a) Mask details are better preserved.
(b) Texture details are enhanced by Multi-modal DreamBooth.
Figure 7: Results for mask processing and texture enhancement.
Deferred point cloud rendering and training. Here we adopt a process akin to that described in Make-it-3D [64] for point cloud rendering and training, but we incorporate our subject-specific prior. It is worth to mention that the point cloud building and texture projection is different, which we add an illustration figure in the appendix for better clarity.
Dataset Metrics RealFusion Make-it-3D Magic123 DreamGS Ours
RealFusion CLIP↑ 0.73 0.83 0.82 0.77 0.90
PSNR↑ 16.87 20.01 19.50 18.94 20.50
LPIPS↓ 0.193 0.119 0.100 0.111 0.094
Test-data CLIP↑ 0.70 0.79 0.86 0.79 0.88
PSNR↑ 16.41 19.31 20.10 19.80 20.30
LPIPS↓ 0.197 0.121 0.098 0.103 0.096
TABLE I: Comparison of different methods on RealFusion15 and Test-data. We compute LPIPS and PSNR under the reference view, and CLIP- Score under novel views.
[…]
TABLE II: The quantitative ablation study result.
IV Experiments
IV-A Comparisons with the State of the Arts
Baselines. We compare our approach against recent state-of-the-art methods: RealFusion [40], Make-It-3D [64], Magic123  [49] and DreamGaussian [63]. We compare these methods using the Realfusion dataset [40] and a customized dataset we created. The Realfusion dataset includes many natural images, while our customized dataset comprises real images and images generated by Stable Diffusion XL [47].
Qualitative comparison. We present a comprehensive collection of qualitative results in Figure 3. RealFusion often generates flat 3D results with colors and shapes that diverge significantly from the input image. Make-it-3D exhibits competitive texture quality but suffers from an issue known as long geometry in side views, particularly noticeable in the reconstruction of objects such as chairs. Magic123 produces visually plausible structures but grapples with a notable issue of multi-face, as it tends to replicate the reference texture in the back view. The geometry generated by DreamGaussian seems to be highly disordered. Under novel views, the newly generated textures are also of poor quality, lacking in detail.
To further demonstrate the effectiveness of the proposed multi-modal DreamBooth, we present results from the coarse stage without any components in the refine stage in Figure 4 compared to Magic123[49]. It can be seen that our results achieve superior results in both geometry and texture, e.g., ours correctly generate the back view of anime girl while Magic123 suffers from multi-face issue.
Quantitative comparison. We quantitatively compare with the baselines in Table I. We use metrics following [64, 49] with PSNR, LPIPS [78], and CLIP-similarity [50]. As shown in the table, we achieves best performance.
V Conclusions and Discussions
Subject-specific prior: multi-modal DreamBooth. We first ablate the effect of using the proposed multi-modal DreamBooth in Figure 5. We observe a consistent improvement in terms of both texture and geometry. Take banana for instance, without multi-modal DreamBooth, the method fails to reconstruct slender structures already visible in the input banana image, which also appears to have unsuccessful reconstruction in Make-it-3D such as the armchair leg shown in Fig 3. The reason is inconsistency stemmed from SDS loss (novel view) and reconstruction loss (reference view). To achieve a balance between these , Magic123 addresses this with additional 3D prior[34], and we alleviates it through multi-modal DreamBooth. This example indeed demonstrate the efficacy of our proposed multi-modal DreamBooth. Also, we conduct a quantitative study on our test benchmark, validating the efficacy of our approach in Table II.
Shading-mode-aware guidance. We ablate shading-mode guidance and found that with it, our training in the coarse stage converged more easily, showing better performance in both geometry and texture, as illustrated in second column and fourth column of Figure 6. This fully demonstrates that shading-mode guidance effectively utilizes the subject’s multi-modal information to provide more robust supervision.
Texture and mask enhancement. Here we ablate the proposed texture and mask enhancement procedures. Figure 7(b) shows that the texture details can be significantly improved by utilizing our multi-modal DreamBooth model. As shown in Figure 7(a), our method can better retain the geometry.
VI Acknowledgement
This work was supported by the National Natural Science Foundation of China (62476011).